MOA-II Obsevation Log Monthly Summary 2006/9

Return to index

Previous month Next month

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB3943-44745252853133060
LMC597-64749527101100
SMC186-2314436171700
ALERT-000
GRB-000
TOTAL---6183741305860 (84h57m40s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
97107979242487107934398711396
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
8983849484801111221211350


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
353332292935464443423637
Field13Field14Field15Field16
424200


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
434242423939383838000


SPECIAL TARGET NAME

CategoryTargetNrun


CCD TEMPERATURE Previous month Next month




VACUUM Previous month Next month




TELESCOPE TEMPERATURE Previous month Next month




COMMENTS
2006/9/1: No observation because of the strong wind all night long. / night time = 39720(sec)

2006/9/2: D1369(30sec) and D1371(300sec) are OK. The wind-speed meter was broken cause of the strong wind. / night time = 39600(sec)

2006/9/3: D1374(30sec), D1375(200sec) are OK. / night time = 39420(sec)

2006/9/4: No observing / night time = 39240(sec)

2006/9/5: D1377 (30sec) , D1379 (100sec) OK. / night time = 39060(sec)

2006/9/6: DARK D1384 (30 sec), D1386 (100 sec) OK. / night time = 38880(sec)

2006/9/7: dark D1392(30sec), D1393(100sec) are OK. Full moon. / night time = 38760(sec)

2006/9/8: dark D1399(30sec), D1400(100sec) are OK. CCD-temperature sensor was broken, and now under repair. / night time = 38580(sec)

2006/9/9: / night time = 38340(sec)

2006/9/10: DARK D1423(60sec), D1425(30sec), D1426(150sec) are OK. / night time = 38160(sec)

2006/9/11: DARK : D1430,1431 (0 sec), D1432 (30 sec), D1433 (60 sec), D1434 (100 sec), D1435 (200 sec) are OK. / night time = 37980(sec)

2006/9/12: DARK : D1439 (60 sec), D1441 (200 sec) are OK. / night time = 37800(sec)

2006/9/13: DARK : D1451 (60 sec), D1454 (300 sec) are OK. / night time = 37620(sec)

2006/9/14: DARK : D1468 (300 sec), D1469 (60 sec) , Today two accidents happened. Vaccum sensor and GPS sensor were error. / night time = 37500(sec)

2006/9/15: DARK : D1482 (60 sec) is OK.. Tonight it was windy(silghtly cloudy or rainy). / night time = 37260(sec)

2006/9/16: We took DARK images every five degrees of CCD temperature because of strong wind. / night time = 37080(sec)

2006/9/17: It was windy all night.And we could not take image because of the GPS error. / night time = 36900(sec)

2006/9/18: We could not take image because of GPS error. / night time = 36720(sec)

2006/9/19: When we were taking L633-lmc12,GPS error appered on MOATEL display,so stopped observation.When we restarted observation,the error disappered. DARK: D1523(60sec) and D1524(300sec) are OK. / night time = 36480(sec)

2006/9/20: We didn't observe because of strong wind. / night time = 36300(sec)

2006/9/21: DARK : D1531 (60 sec), D1532 (300 sec) are OK. / night time = 36120(sec)

2006/9/22: DARK: D1537 (60 sec) , D1538 (300 sec) Now, data has pointing erroor (remain unresoved). / night time = 35940(sec)

2006/9/23: It was windy or rainy tonight. / night time = 35700(sec)

2006/9/24: / night time = 35520(sec)

2006/9/25: D1539(60sec) is OK. We observed GB. / night time = 35340(sec)

2006/9/26: DARK: D1542 (60 sec), D1544 (300 sec) are OK. / night time = 35100(sec)

2006/9/27: DARK: D1548 (60 sec) and D1552(300 sec) are OK. / night time = 34920(sec)

2006/9/28: DARK: D1554 (60 sec), D1555 (300 sec) / night time = 34740(sec)

2006/9/29: DARK:D1558(60sec), D1566(40sec), D1569(200sec) and D1572(300sec) are OK. GPS error happend at 10:45pm. / night time = 34440(sec)

2006/9/30: DARK:D1578(60sec) and D1581(300sec) are OK. / night time = 37860(sec)

TOTAL NIGHT TIME: 1117080(sec)

OPERATING RATE: 40.8 %
GB: 22.1 %
LMC: 10.9 %
SMC: 7.7 %
ALERT: 0.0 %
GRB: 0.0 %

Return to index

Previous month Next month